# GPT-4V with Emotion: A Zero-shot Benchmark for Generalized Emotion Recognition

### 03 Feb 2024 v2

@@ Abstract
Recently, GPT-4 with Vision (GPT-4V) has demonstrated remarkable visual capabilities across various tasks, but its performance in emotion recognition has not been fully evaluated. To bridge this gap, we present the quantitative evaluation results of GPT-4V on 19 benchmark datasets covering 5 tasks: visual sentiment analysis, micro-expression recognition, facial emotion recognition, dynamic facial emotion recognition, and multimodal emotion recognition. This paper collectively refers to these tasks as “Generalized Emotion Recognition (GER)”. Through experimental analysis, we observe that GPT-4V generally outperforms supervised systems in visual sentiment analysis, highlighting its powerful visual understanding capabilities. Meanwhile, GPT-4V shows the ability to integrate multimodal clues and exploit temporal information, which is also critical for emotion recognition. Despite these achievements, GPT-4V is primarily tailored for general-purpose domains, which cannot recognize micro-expressions that require specialized knowledge. To the best of our knowledge, this paper provides the first quantitative assessment of GPT-4V for the GER tasks, offering valuable insights to researchers in this field. It can also serve as a zero-shot benchmark for subsequent research. Our code and evaluation results are available at: https://github.com/zeroQiaoba/gpt4v-emotion.
@@ 1 Introduction
Emotion has attracted increasing attention from researchers due to its importance in human-computer interaction cowie2001emotion ; brave2007emotion . Current research on emotion recognition primarily centers on two aspects: one aims to identify the emotions evoked by stimuli, essentially predicting how viewers might feel after seeing these stimuli you2017visual ; zhao2021affective ; the other aims to analyze human emotions in images li2019blended ; li2020deep or videos lian2022smin ; lian2024merbench . This paper collectively refers to these tasks as “Generalized Emotion Recognition (GER)”.
@@
Emotion can be conveyed through various modalities, including text, audio, and video picard2000affective . Among them, visual information (such as colorfulness, brightness, facial expression, human action, etc) contains rich emotion-related content yang2023emoset and plays a crucial role in GER tasks. To enhance visual understanding capabilities, researchers have proposed various algorithms and achieved noteworthy progress poria2017review . With the development of deep learning, current research on GER has shifted from handicrafted features to deep neural networks. Recently, GPT-4V openai2023gpt4v has showcased impressive visual understanding capabilities across various tasks. This leads to a question: can GPT-4V solve the GER problem to some extent? If so, what is the future direction following the emergence of GPT-4V?
@@
In September 2023, GPT-4V was integrated into ChatGPT openai2022chatgpt , triggering a series of user reports aimed at investigating its visual capabilities yang2023dawn ; lyu2023gpt . However, these reports generally involve a limited number of samples per task and can only provide qualitative insights into GPT-4V. Although OpenAI released the API in November 2023, it was initially limited to 100 requests per day. Therefore, it remains difficult to evaluate GPT-4V against state-of-the-art systems on benchmark datasets. Recently, OpenAI has increased the daily limit, allowing us to conduct more comprehensive evaluations.
@@
In this paper, we provide quantitative evaluation results of GPT-4V on GER tasks, covering visual sentiment analysis ortis2020survey , micro-expression recognition li2022deep , facial emotion recognition li2020deep , dynamic facial emotion recognition wang2022ferv39k , and multimodal emotion recognition lian2024merbench . Figure 1 shows the overall results of GPT-4V. We also report results of random guessing (a heuristic baseline that randomly selects labels from candidate categories) and supervised systems. To fairly compare different methods, we conduct experiments on benchmark datasets and use the same evaluation metric. As shown in Figure 1, overall, GPT-4V outperforms random guessing but still lags behind supervised systems. To figure out the reason, we conduct a comprehensive analysis of GPT-4V’s multi-faceted performance, including multimodal fusion ability, temporal modeling capability, color space robustness, prediction consistency, etc (see Section 5). We hope this paper can inspire subsequent researchers and shed light on tasks that GPT-4V can effectively address and those that require further exploration. The main contribution of this paper can be summarized as follows:
@@
    To the best of our knowledge, this is the first work to quantitatively evaluate the performance of GPT-4V in emotion recognition. Additionally, this paper can serve as a zero-shot benchmark for subsequent research.
@@
    Our evaluation results demonstrate that GPT-4V focuses on general-purpose domains. For micro-emotion recognition that needs specialized knowledge, it performs even worse than random guessing in some datasets. Meanwhile, we observe that GPT-4V can integrate multimodal clues and capture temporal information during inference.
@@
    This paper can offer guidance for future directions following the emergence of GPT-4V. We open-source the code and encourage subsequent researchers to broaden the evaluation scope by including more tasks and datasets.
The remainder of this paper is organized as follows: In Section 2, we briefly review recent works. In Sections 3 and 4, we provide a comprehensive overview of GER tasks and datasets, along with a detailed description of the GPT-4V calling strategy. In Section 5, we report results and conduct an in-depth analysis of GPT-4V’s multi-faceted capabilities. Finally, we conclude this paper and discuss future work in Section 6.
Figure 1: Performance of different methods on GER. “Random” is a heuristic baseline that randomly selects labels from candidate categories.
@@ 2.1 Generalized Emotion Recognition
GER aims to identify the emotions of individuals or to recognize the emotions that individuals might evoke from stimuli. It covers a variety of tasks, and this paper focuses on five representative GER tasks. In this section, we detail the similarities and distinctions between these tasks.
@@
Visual sentiment analysis ortis2020survey investigates the impact of image stimuli on people’s emotions, where the image content is not necessarily human-centric. In contrast, facial emotion recognition li2020deep aims to identify emotions in human-centric images, particularly focusing on observable macro-expressions. In addition to macro-expressions, there is another kind of expression called micro-expressions, which are characterized by short duration, low intensity, and sparse facial action units. The task of identifying such expressions is called micro-expression recognition li2022deep . All of the aforementioned tasks identify emotions from static images. Dynamic facial emotion recognition wang2022ferv39k extends the analysis from static images to image sequences, requiring the use of temporal information. Moreover, emotions can be conveyed through modalities other than visuals, such as text and audio. Therefore, multimodal emotion recognition lian2024merbench is introduced to integrate multimodal information for a more comprehensive understanding of emotions.
2.2 Multimodal Large Language Model
Recently, large language models (LLMs) with massive parameters have attracted increasing attention due to their unprecedented performance penedo2023refinedweb ; touvron2023llama ; workshop2022bloom . In GER tasks, visual information plays an important role in emotion recognition, but LLMs only support lexical inputs. Therefore, this section focuses on multimodal LLMs that support visual inputs.
Among all multimodal LLMs, MiniGPT-4 zhu2023minigpt and LLaVA liu2023visual are two representative models. They enhance the visual capability by aligning the frozen visual encoder with LLMs using a simple projection layer., which preserves most of the information encoded in pre-trained models and achieves good performance in visual understanding. This approach is further extended by other works and applies to support more modalities, such as video li2023videochat , audio zhang2023speechgpt , and depth images su2023pandagpt . Currently, GPT-4V openai2023gpt4v is the state-of-the-art multimodal LLM for various tasks alpaca_eval ; lu2024mathvista . However, its performance in emotion recognition has not been fully studied. Hence, this paper evaluates the performance of GPT-4V on GER, providing a valuable zero-shot benchmark for other multimodal LLMs.
2.3 Evaluation on GPT-4V
With the emergence of GPT-4V, a large number of user reports have surfaced to evaluate its visual capabilities lin2023mm ; zhou2023exploring ; wu2023can . Among them, Yang et al. yang2023dawn conducted the most comprehensive assessment, covering coding, captioning, and reasoning abilities. However, they primarily performed qualitative analysis on limited samples and did not disclose the performance gap between GPT-4V and supervised systems on benchmark datasets. To bridge this gap, there have been efforts to quantitatively analyze GPT-4V’s mathematical reasoning lu2024mathvista and visual abilities wu2023gpt4vis , but none have specifically explored its emotional understanding capabilities. In this paper, we provide the first quantitative evaluation of GPT-4V on GER tasks. It is worth noting that, unlike other tasks, emotions are related to humans. However, human identity is treated as sensitive information in GPT-4V, which raises new requirements for the calling strategy. More details can be found in Section 4.
@@ 2.3 Evaluation on GPT-4V
Micro-expression Rec. (Image) CASME yan2013casme 195 ACC happiness, sadness, fear, surprise, disgust, repression, contempt, tense
@@
Facial Emotion Rec. (Image) CK+ lucey2010extended 981 ACC happiness, sadness, anger, fear, disgust, surprise, contempt
@@
Dynamic Facial Emotion Rec. (Video) DFEW (fd1) jiang2020dfew 2,341 WAR happiness, sadness, anger, fear, disgust, surprise, neutral
@@
Multimodal Emotion Rec. (Video, Audio, Text) CMU-MOSI zadeh2017tensor 686 WAF sentiment intensity
@@ 3 Task Description
In this paper, we conduct a comprehensive evaluation across 19 datasets covering 5 tasks. Table 1 summarizes statistics of different datasets. To ensure a fair comparison with supervised systems, we evaluate performance on the official test set and select the most common evaluation metric. Figure 2 shows examples of different datasets, demonstrating the diversity of dataset types. Some datasets are in-the-wild (such as AffectNet), while others are laboratory-controlled (such as CASME and CK+). Meanwhile, there is a disparity in color spaces across distinct datasets. Some datasets use grayscale images (such as CK+), while others use RGB images (such as CASME and AffectNet). This section describes each task and dataset in detail.
@@
Facial Emotion Recognition chooses five benchmark datasets: CK+ lucey2010extended , FERPlus barsoum2016training , SFEW 2.0 dhall2015video , RAF-DB li2017reliable , and AffectNet mollahosseini2017affectnet . The first two datasets contain grayscale images, while the last three datasets contain RGB images. Specifically, CK+ lucey2010extended includes 593 video sequences from 123 subjects. Following previous work jiang2022disentangling , we extract the last three frames of each sequence to construct the dataset. FERPlus barsoum2016training is an extension of FER2013 goodfellow2013challenges , where each sample is relabeled by ten annotators. SFEW 2.0 dhall2015video extracts keyframes from movie clips, encompassing diverse head poses, occlusions, and illuminations. RAF-DB li2017reliable contains thousands of samples with basic and compound expressions, and this paper focuses on basic emotions. AffectNet mollahosseini2017affectnet has 8 labels and each label contains 500 samples for evaluation.
Table 2: Prompt template for different tasks. Please replace “Candidate labels” in the following prompts into labels of each dataset. This paper uses batch-wise calling strategy to handle request limits. By default, we set the batch size to 20 for image-level inputs and 6 for video-level inputs.
@@
Micro-expression Rec.
@@
Dynamic Facial Emotion Recognition focuses on more challenging image sequences. In this task, we conduct experiments on four benchmark datasets: FERV39k wang2022ferv39k , RAVDESS livingstone2018ryerson , eNTERFACE05 martin2006enterface , and DFEW jiang2020dfew . For the first three datasets, we use the official training/validation/test splits and evaluate the performance on the official test set. DFEW jiang2020dfew contains 5 folds with a total of 11,697 samples. To reduce the evaluation cost, we only report the results of fold 1 (fd1), consistent with previous works wang2022dpcnet ; chen2023static .
Multimodal Emotion Recognition aims to integrate various modalities (including audio, video, and text) to identify emotions. For this task, we select three benchmark datasets: CH-SIMS zadeh2017tensor , CMU-MOSI yu2020ch , and MER-MULTI lian2023mer . The first two datasets offer sentiment intensity scores for each sample. In our evaluation, we concentrate on the negative/positive classification task, where positive and negative classes are assigned for $<0$ and $>0$ scores, respectively. MER-MULTI is a subset of the MER2023 dataset lian2023mer , providing both discrete and dimensional labels for each sample. In this paper, we focus on the recognition performance of discrete emotions.
4 GPT-4V Calling Strategy
This paper evaluates the performance of the latest GPT-4V API, “gpt-4-vision-preview”. GER tasks involve diverse modalities, such as image, text, video, and audio. However, the current version of GPT-4V has limitations and only supports image and text inputs. To process video data, we sample the video and convert it into multiple images. For audio data, we try converting it to a mel-spectrogram, but GPT-4V fails to generate proper responses for this input. Therefore, the primary focus of this paper revolves around image, text, and video. To meet the request limits of API and reduce rejected cases caused by security checks, we propose batch-wise calling and recursive calling strategies. More details are provided below.

### 18 Mar 2024 v3

@@ Abstract
Recently, GPT-4 with Vision (GPT-4V) has demonstrated remarkable visual capabilities across various tasks, but its performance in emotion recognition has not been fully evaluated. To bridge this gap, we present the quantitative evaluation results of GPT-4V on 21 benchmark datasets covering 6 tasks: visual sentiment analysis, tweet sentiment analysis, micro-expression recognition, facial emotion recognition, dynamic facial emotion recognition, and multimodal emotion recognition. This paper collectively refers to these tasks as “Generalized Emotion Recognition (GER)”. Through experimental analysis, we observe that GPT-4V exhibits strong visual understanding capabilities in GER tasks. Meanwhile, GPT-4V shows the ability to integrate multimodal clues and exploit temporal information, which is also critical for emotion recognition. However, it’s worth noting that GPT-4V is primarily designed for general domains and cannot recognize micro-expressions that require specialized knowledge. To the best of our knowledge, this paper provides the first quantitative assessment of GPT-4V for GER tasks. We have open-sourced the code and encourage subsequent researchers to broaden the evaluation scope by including more tasks and datasets. Our code and evaluation results are available at: https://github.com/zeroQiaoba/gpt4v-emotion.
@@ 1 Introduction
Emotion has attracted increasing attention from researchers due to its importance in human-computer interaction. Current research on emotion recognition primarily centers on two aspects: one aims to identify the emotions evoked by stimuli, essentially predicting how viewers might feel after seeing these stimuli you2017visual ; the other aims to analyze the emotions conveyed by humans in various ways lian2022smin . We collectively refer to these tasks as “Generalized Emotion Recognition (GER)”.
@@
Emotions are associated with lexical, visual, and acoustic information. Among them, visual information (such as colorfulness, brightness, facial expression, human action, etc) contains rich emotion-related content yang2023emoset . To enhance visual understanding capabilities, researchers have proposed various algorithms and achieved noteworthy progress. With the development of deep learning, current research has shifted from handcrafted features to deep neural networks. Recently, GPT-4V has showcased impressive visual understanding capabilities across various tasks and domains. This leads to a question: can GPT-4V solve the GER problem to some extent? If so, what is the future direction following the emergence of GPT-4V?
@@
In September 2023, GPT-4V was integrated into ChatGPT, triggering a series of user reports aimed at investigating its visual capabilities yang2023dawn . At that time, OpenAI had not yet released the API service, and users could only manually upload test samples to the web service. Due to the high manual effort required, these reports generally involved a limited number of samples per task and merely provided qualitative insights into GPT-4V. In November 2023, OpenAI released the API, but it was initially limited to 100 requests per day. It remained difficult to evaluate GPT-4V against state-of-the-art systems on benchmark datasets. Recently, OpenAI has increased the daily limit, allowing us to conduct more comprehensive evaluations.
@@
In this paper, we provide quantitative evaluation results of GPT-4V on GER tasks, covering visual sentiment analysis ortis2020survey , tweet sentiment analysis niu2016sentiment , micro-expression recognition li2022deep , facial emotion recognition li2020deep , dynamic facial emotion recognition wang2022ferv39k , and multimodal emotion recognition lian2024merbench . Figure 1 shows the overall results of GPT-4V. We also report results of random guessing (a baseline that randomly selects labels from candidate classes) and supervised systems. To fairly compare different methods, we conduct experiments on benchmark datasets and use the same evaluation metric. Overall, GPT-4V outperforms random guessing but still lags behind supervised systems. To figure out the reason, we further conduct a comprehensive analysis of GPT-4V’s multi-faceted capabilities, including multimodal fusion, temporal modeling, robustness, stability, etc. We hope this paper can inspire subsequent researchers and shed light on tasks that GPT-4V can effectively address and those that require further exploration. The main contribution of this paper can be summarized as follows:
@@
    To the best of our knowledge, this is the first work to quantitatively assess GPT-4V’s performance in GER tasks.
@@
    Our evaluation results indicate that GPT-4V is primarily designed for general-purpose domains. It performs poorly in micro-expression recognition that requires specialized knowledge. Meanwhile, it can integrate multimodal clues and capture temporal information during inference.
@@
    This paper can provide guidance on potential future directions following the emergence of GPT-4V. It can also serve as a zero-shot benchmark for subsequent research.
The remainder of this paper is organized as follows: In Section 2, we briefly review recent works. In Sections 3 and 4, we provide an overview of GER tasks and datasets, along with a detailed description of our GPT-4V calling strategy. In Section 5, we report results and conduct an in-depth analysis of GPT-4V’s multi-faceted capabilities. In Sections 6 and 7, we summarize the main challenges and limitations of GPT-4, discuss potential directions for future research, and conclude the entire paper.
Figure 1: Performance of different methods on GER tasks. Here, “random” is a heuristic baseline that randomly selects labels from candidate categories.
@@ 2.1 Generalized Emotion Recognition
GER covers a variety of tasks and this paper focuses on 6 representative tasks. In this section, we detail the similarities and distinctions among them. Visual sentiment analysis ortis2020survey explores the impact of image stimuli on people’s emotions, while tweet sentiment analysis niu2016sentiment aims to determine how individuals feel when posting visual and textual content on social networks. In these tasks, the image is not necessarily human-centric. In contrast, facial emotion recognition li2020deep aims to identify emotions in human-centric images, particularly focusing on observable macro-expressions. Besides macro-expressions, there is another type of expression known as micro-expressions, which are characterized by short duration, low intensity, and sparse facial action units. The task of identifying such expressions is termed micro-expression recognition li2022deep . All of the aforementioned tasks identify emotions from static images. Dynamic facial emotion recognition wang2022ferv39k extends the analysis from static images to image sequences, requiring the use of temporal information. Furthermore, emotions can be conveyed through multiple modalities. Therefore, multimodal emotion recognition lian2024merbench is introduced to integrate multimodal clues for a more comprehensive understanding of emotions.
2.2 Unimodal Emotion Recognition
Emotions can be conveyed through multiple modalities, such as images, text, and audio. Since the early 1970s, Ekman et al. ekman1978facial ; ekman1994strong have conducted prior studies on facial expressions and developed the Facial Action Coding System (FACS) to code facial expressions. In this system, the face can be described as a set of action units, each associated with a muscular basis. Besides facial expressions, emotions can also be conveyed through text and audio, which carry important communicative messages. Text and audio are highly correlated. On the one hand, we can recognize the text from audio. On the other hand, if we ignore the manner in which the text is spoken, we might misunderstand the meaning of the text sebe2005multimodal . With the development of deep learning, researchers have proposed various effective frameworks for emotion recognition lian2021ctnet . These models are usually task-specific, i.e., we need to train on emotion corpora. In contrast, this paper uses GPT-4V, which can recognize emotions in a zero-shot manner without further training.
2.3 Multimodal Emotion Recognition
Humans express emotions in a complementary manner across various modalities. To better understand emotions, we need to integrate multimodal clues. According to the fusion location, current methods can be roughly divided into three categories: feature-level fusion, model-level fusion, and decision-level fusion wu2014survey . Specifically, feature-level fusion concatenates unimodal features at the input level, while decision-level fusion combines unimodal decision results. Despite their promising performance, these methods ignore interactions and correlations between different modalities. Therefore, researchers propose model-level fusion to capture multimodal dependencies. Recently, Lian et al. lian2024merbench built MERBench and conducted a systematic analysis of different model-level fusion strategies. Unlike the above methods, GPT-4V achieves multimodal fusion through prompts. More details can be found in Table 2.
@@ 2.3 Multimodal Emotion Recognition
Tweet Sentiment Analysis (Image, Text) MVSA-Single niu2016sentiment 451 ACC positive, neutral, negative
MVSA-Multiple niu2016sentiment 1,702 ACC positive, neutral, negative
Micro-expression Recognition (Image) CASME yan2013casme 195 ACC happiness, sadness, fear, surprise, disgust, repression, contempt, tense
@@
Facial Emotion Recognition (Image) CK+ lucey2010extended 981 ACC happiness, sadness, anger, fear, disgust, surprise, contempt
@@
Dynamic Facial Emotion Recognition (Video) DFEW (fd1) jiang2020dfew 2,341 WAR happiness, sadness, anger, fear, disgust, surprise, neutral
@@
Multimodal Emotion Recognition (Video, Audio, Text) CMU-MOSI zadeh2017tensor 686 WAF sentiment intensity
@@
2.4 Multimodal Large Language Model
Large language models (LLMs) with massive parameters have attracted increasing attention due to their unprecedented performance. In GER tasks, visual information plays an important role in emotion recognition, but LLMs only support lexical inputs. Therefore, we focus on multimodal LLMs (MLLMs) that support visual inputs. Among all MLLMs, MiniGPT-4 zhu2023minigpt and LLaVA liu2023visual are two representative models. They enhance the visual capability by aligning the frozen visual encoder with LLMs using a simple projection layer, which preserves most of the information encoded in pre-trained models and achieves good performance in visual understanding. This approach is further extended by other works and applies to support more modalities, such as video li2023videochat , audio zhang2023speechgpt , and depth images su2023pandagpt . Currently, GPT-4V is the state-of-the-art MLLM for various tasks alpaca_eval ; lu2024mathvista . However, its performance in emotion recognition has not been fully studied. Hence, this paper evaluates the performance of GPT-4V on GER tasks, providing a valuable zero-shot benchmark for other MLLMs.
2.5 Evaluation on GPT-4V
With the emergence of GPT-4V, a large number of user reports have surfaced to evaluate its visual ability lin2023mm ; wu2023can . Among them, Yang et al. yang2023dawn conducted the most comprehensive assessment, covering coding, captioning, and reasoning abilities. However, they primarily performed qualitative analysis on limited data and did not disclose the performance gap between GPT-4V and supervised systems on benchmark datasets. To bridge this gap, there have been efforts to quantitatively analyze GPT-4V’s mathematical reasoning lu2024mathvista and visual abilities wu2023gpt4vis , but none have specifically explored its emotional understanding capabilities. In this paper, we provide the first quantitative evaluation of GPT-4V on GER tasks. It is worth noting that the emotion is related to humans. However, human identity is treated as sensitive information in GPT-4V, which raises new requirements for the calling strategy. To this end, we also design a calling strategy specifically for GER tasks.
@@ 3 Task Description
In this section, we describe each task and dataset in detail. Table 1 summarizes statistics of different datasets. To ensure a fair comparison with supervised systems, we evaluate performance on the official test set and select the most common evaluation metric. Figure 2 shows examples of different datasets, demonstrating the diversity of dataset types. Some datasets are in-the-wild (such as AffectNet), while others are laboratory-controlled (such as CASME and CK+). Meanwhile, there is a disparity in color spaces across distinct datasets. Some datasets use grayscale images (such as CK+), while others use RGB images (such as CASME and AffectNet).
@@
Tweet Sentiment Analysis aims to determine how people feel when posting visual and textual content on social networks. We conduct experiments on two benchmark datasets: MVSA-Single niu2016sentiment and MVSA-Multiple niu2016sentiment . These datasets contain image-text pairs collected from Twitter, and annotators assign one of three sentiments (positive, negative, and neutral) to the image and text respectively. MVSA-Single contains 5,129 image-text pairs labeled by a single annotator. MVSA-Multiple is an extended version of MVSA-Single with more samples and more annotators. For a fair comparison, we adopt the same dataset processing strategy in previous works, and then randomly select 10% of the samples as the test set xu2017multisentinet ; zhu2022multimodal .
Facial Emotion Recognition focuses on five benchmark datasets: CK+ lucey2010extended , FERPlus barsoum2016training , SFEW 2.0 dhall2015video , RAF-DB li2017reliable , and AffectNet mollahosseini2017affectnet . The first two datasets contain grayscale images, while the last three datasets contain RGB images. Specifically, CK+ lucey2010extended includes 593 video sequences from 123 subjects. Following previous work jiang2022disentangling , we extract the last three frames of each sequence to construct the dataset. RAF-DB li2017reliable contains thousands of samples with basic and compound expressions, and this paper focuses on basic emotions. AffectNet mollahosseini2017affectnet has 8 labels and each label contains 500 samples for evaluation. For FERPlus barsoum2016training and SFEW 2.0 dhall2015video , we use the official test set for performance evaluation.
Table 2: Prompt template for different tasks. Please replace “Candidate labels” in the following prompts into candidate labels of each dataset. This paper uses batch-wise calling strategy to handle request limits. By default, we set the batch size to 20 for image-level inputs and 6 for video-level inputs.
@@
Tweet Sentiment Analysis
Please play the role of an emotion recognition expert. We provide 20 image-text pairs. Please analyze how he will feel if he post this image-text pair on social media. If there is a person in the image, ignore that person’s identity. For each image-text pair, please sort the provided categories from high to low according to the similarity with the input image-text pair. Here are the optional categories: {Candidate labels}. The output format should be {’name’:, ’result’:} for each image-text pair.
Micro-expression Rec
@@
Dynamic Facial Emotion Recognition focuses on more challenging image sequences. In this task, we conduct experiments on four benchmark datasets: FERV39k wang2022ferv39k , RAVDESS livingstone2018ryerson , eNTERFACE05 martin2006enterface , and DFEW jiang2020dfew . For the first three datasets, we evaluate the performance on the official test set. DFEW jiang2020dfew contains 5 folds with a total of 11,697 samples. To reduce the evaluation cost, we only report the results of fold 1 (fd1), consistent with previous works wang2022dpcnet ; chen2023static .
Multimodal Emotion Recognition focuses on three benchmark datasets: CH-SIMS zadeh2017tensor , CMU-MOSI yu2020ch , and MER-MULTI lian2023mer . The first two datasets offer sentiment intensity scores. In this paper, we concentrate on the negative/positive classification task, where positive and negative classes are assigned for $<0$ and $>0$ scores, respectively. MER-MULTI is a subset of the MER2023 dataset lian2023mer , providing both discrete and dimensional labels for each sample. In this paper, we focus on the recognition performance of discrete emotions.